# Solving High-dimensional Inverse Problems Using Amortized Likelihood-free Inference with Noisy and Incomplete Data

### v1

@@ 3.5 Application
As described in Section 2.1, our case study focuses on estimating the hydraulic log-conductivity field of the 706 control volumes. The observed data consists of synthetic, noisy observations of the hydraulic head at 13 observation locations over 25 timesteps. The log-conductivity prior distribution is a Gaussian process prior with zero mean and the exponential covariance function. From this prior, a total of 5,001 synthetic datasets are generated using MODFLOW. Fig. 7 illustrates a single realization of this synthetic data, showing the input conductivity field alongside the hydraulic head field across the simulation domain at the 25 timesteps. Cells colored white are inactive, meaning the governing equation is not solved in these cells, while the remaining cells are active. Additionally, to simulate measurement errors, random Gaussian noise with zero mean and standard deviation of 0.01 is added to the synthetic observations.
Figure 7: A realization synthetic data at 706 locations: (a) log-conductivity field; (b) hydraulic head field
@@ 3.5.1 Model training
Fig. 8 shows the performance of the trained model for 100 test datasets. Due to space limitations in the paper, we show results for only a subset of the 706 parameters. The error bars depict the $95\%$ credibility intervals, calculated from 2,000 posterior samples. The results show that the estimated parameters generally align well with the ground truth, as evidenced by 87% (13 out of 15) of the selected parameters achieving a achieving a coefficient of determination $R^{2}$ exceeding 0.8. Furthermore, the correlation coefficient $(r)$ confirms a strong linear relationship between the estimated posterior mean and actual values. However, it is noted that certain parameters, such as $\theta_{312}$ and $\theta_{315}$, exhibit lower $R^{2}$ values compared to others. This variation is attributed to their relative insensitivity to the observations collected from the 13 sensor locations. The 15 parameters shown were selected randomly for illustrative purposes.
Figure 8: Validation of trained model performance
@@ 3.5.2 Results of parameter estimation
We compare our method against PEST-IES for two test problems corresponding to two reference conductivity fields. It is important to highlight the workflow differences between the methods: PEST-IES is likelihood-based and so relies on a forward model to estimate parameters, whereas our proposed method bypasses the evaluation of the likelihood function. We employed three metrics to compare the accuracy of the parameter estimates derived from both methods:
@@
    where $\theta_{i,\text{est}}$ and $\theta_{i,\text{ref}}$ are the $i$-th estimated and true parameters, respectively. A lower relative error signifies higher accuracy in the parameter estimation.
@@
    A higher LPP value points to a more accurate model estimation, as it reflects a better fit of the estimated parameters to the true values.
Figs. 9 and 10 display the posterior mean log-conductivity fields estimated with PEST-IT and the likelihood-free method, true log-conductivity fields, and point relative errors for tests #1 and #2, respectively. It can be seen that both methods yield comparably accurate estimations, which suggests that the proposed method can reliably reconstruct the conductivity field without the need for likelihood function evaluation. Figs. 11 and 12 show the true parameter values, the posterior mean values, and the corresponding $95\%$ credibility intervals (shown in grey shaded areas) for both PEST-IES and the proposed method. A majority of the estimated parameters are contained within the credibility interval, which suggests that the proposed method is reliable and yields estimations that are comparable to those obtained via the PEST-IES benchmark method. Furthermore, these figures highlight that parameter estimation for test #2 poses a greater challenge, evidenced by the larger variability among the 706 parameters, particularly in the 200-300 and 500-600 location ranges, in comparison to test #1. Despite these challenges, the proposed method manages to produce accurate parameter estimations, demonstrating its efficacy even under challenging conditions.
Figure 9: Estimation of log-conductivity field using test #1: (a) PEST-IES (b) Proposed
Figure 10: Estimation of log-conductivity field using test #2: (a) PEST-IES (b) Proposed
Figure 11: Estimation of 706 parameters using test #1
Figure 12: Estimation of 706 parameters using test #2
Table 1 presents the accuracy metrics for the proposed PEST-IES methods for two test problems. For test #1, both methods exhibit comparable accuracy, as measured by the relative $\ell_{2}$ error and the LPP, with the proposed method showing a marginally lower average relative error compared to PEST-IES. When evaluating test #2, the proposed method demonstrates slightly better performance, with a lower average relative error, a lower relative $\ell_{2}$ error, and a marginally higher LPP score. While these differences are modest, they suggest that the proposed method maintains its precision in parameter estimation, even with the increased variability among the 706 parameters in test #2. Moreover, while the proposed and PEST-IES show similar performance levels in parameter estimation, the computational efficiency between the two is noteworthy (details are presented in Sec. 3.5.4). The PEST-IES method requires repeated evaluations of the forward model, which is computationally intensive. In contrast, the proposed method does not require the forward model evaluations outside the training data generation stage. This distinction means that once the model is adequately trained, the proposed method can deliver fast inference, offering an advantage in computational cost.
Table 1: Comparison between proposed and PEST-IES method
@@
Fig. 13 shows coverage map for each method and each of the two datasets under consideration. In this figure, yellow control volumes indicate that the true value is within the $95\%$ credibility interval, and a red control volume indicates that the true parameter is outside said interval. For test #1, PEST-IES and the proposed methods demonstrate similar coverages, indicating comparable reliability. However, for test #2, which presents a more complex challenge for parameter estimation, both methods yield lower coverage compared to test #1. Despite this, the proposed method achieves a higher coverage value closer to the expected value of $95\%$.
Figure 13: Parameter coverage: (a) test #1; (b) test #2
Some likelihood-free approaches, such as ABC, and likelihood-based methods, including PEST-IES, require resampling or solving minimization problems for each new set of measurements. One of the attractive attributes of the proposed method is that is does not have to be retrained when measurements change. It can perform parameter estimation across different measurement set sizes utilizing a single trained model. To illustrate this adaptability, we considered five scenarios where the number of measurements at each location varies from 20 to 25 with the time increments of one year between the measurements. In the training dataset, the number of observations at each location is 25. Fig. 14 shows the relative $\ell_{2}$ errors for various numbers of measurements. As anticipated, the relative $\ell_{2}$ errors for both test problems generally decrease as the number of observations increases. Importantly, the findings in Fig. 14 confirm that the proposed model, once trained, can estimate parameters from datasets of arbitrary sizes.
Figure 14: Parameter estimation by varied data length: (a) dataset #1; (b) dataset #2
@@ 3.5.3 Observation prediction
In this section, we evaluate the posterior predictive distribution of hydraulic head measurements obtained using PEST-IES and our proposed method. The posterior predictive distribution is calculated by inputting the estimated log-conductivity field into the MODFLOW model. The proposed method is used to generate 2,000 posterior samples for each parameter, which are fed into the MODFLOW model, resulting in the same number of observation predictions. Figs. 15 and 16 display the predictive observation distribution obtained via PEST-IES and the proposed methods, with the grey shaded areas representing the corresponding $95\%$ credibility intervals. It can be seen that both methods are capable of accurately predicting observations across all observation locations and timesteps. For the PEST-IES method, the predictions closely coincide with the true values for both datasets, with a notably small uncertainty, as indicated by the tight credibility intervals. In contrast, the proposed method exhibits larger uncertainties; the credibility intervals span a wider range than those of PEST-IES, especially for observations at locations #1 through #4. This difference can be attributed to the fact that PEST-IES performs parameter estimation using a forward model that directly generates the synthetic observation data, thus bypassing the need for model training. Consequently, PEST-IES can yield more accurate predictions, albeit with the trade-off of requiring more computational resources due to the repeated evaluation of the forward model.
@@
In the proposed method, the errors due to the summary and inference network training contribute to the parameter estimation error. There are several error sources in the network models. Firstly, the optimization of the loss function inevitably introduces some error due to the finite number of training data, as described in Eq. 15. This error arises from the stochastic nature of the sampling process used during training. Secondly, the summary network choice might be suboptimal, potentially failing to capture all crucial information. Additionally, the designed invertible network might not sufficiently capture the complexities of the forward model’s behavior. For instance, the depth of the inference network in our study, which comprises 10 invertible blocks, may be inadequate for handling the 706-dimensional parameter space. Thirdly, fine-tuning hyperparameters such as the dropout rate, learning rate, and optimization method presents a complex challenge, with limited guidance available for the ideal design and training of neural networks. These errors can propagate into the predictive performance, affecting the uncertainty in observation predictions. Despite these larger uncertainties, the mean predictions from the proposed method consistently align well with the true observations and remain within the confidence intervals, underscoring the method’s reliability. Moreover, for test #2, both the uncertainty and deviation from the true observations are more pronounced than for test #1, as the proposed method finds it more challenging to estimate parameters for test #2—this is evidenced by the higher relative $\ell_{2}$ error reported in Table 1.
@@
Figure 15: Observation prediction at 13 locations for test #1: (a) proposed method; (b) PEST-IES
@@
Figure 16: Observation prediction at 13 locations for test #2: (a) proposed method; (b) PEST-IES
3.5.4 Computational cost
@@
Table 2 provides a comparative analysis of the computational costs associated with the proposed method and PEST-IES for estimating conductivity fields. The proposed method requires approximately 4.5 hours for training, but once this phase is complete, it can perform parameter estimation on a single measurement dataset in about 2 seconds. In contrast, PEST-IES requires around 2.5 hours to complete parameter estimation for each measurement dataset. Although the proposed method demands more time for training, this process can be conducted offline using simulated data. Therefore, the proposed method can be used for real-time parameter estimation.
Table 2: Comparison of computational cost between proposed method and PEST-IES
Proposed PEST-IES
Training (h) Inference (s) Inference (h)
Computational cost 4.5 2 2.5
@@ 4 Conclusion
In this work, we introduce a novel likelihood-free approach using normalizing flow to tackle high-dimensional inverse problems. The proposed method includes three phases. Initially, the first phase involves generating synthetic data from the forward model, utilizing prior knowledge of the model parameters. The second phase focuses on training neural networks. We propose the joint training of two complementary deep neural networks: a summary network and an inference network. The summary network autonomously learns the most informative representation of data, ensuring that the features extracted are maximally beneficial for inference tasks. The inference network is designed using an architecture that combines invertible neural networks with neural spline flows, featuring a stacked configuration of coupled cACLs and cSLs. The summary statistics are input into the inference network, allowing for the joint optimization of the two networks. The final phase is dedicated to real-time inference. Once the model is trained, the posteriors of the model parameters can be computed in real time. We apply this framework to estimate the 706-dimensional hydraulic conductivity field using noisy and incomplete observations from the Freberg groundwater model. Additionally, we conduct a comparative analysis between our method and the likelihood-based benchmark method, PEST-IES. The results highlight that our method can accurately and efficiently estimate the high-dimensional conductivity field. Below, we summarize the main advantages of our proposed method.
First, the proposed method represents a fully likelihood-free approach that directly estimates the posterior without the need to evaluate a likelihood function. Furthermore, it rigorously ensures an efficient sampling process for the true posterior, independent of any assumptions regarding the prior or posterior distributions.
Second, the method exhibits robust performance across varied measurement sequences. In other words, it can handle variable dataset sizes with a single trained model, that is, without the need of re-training for different measurement sequences. This flexibility is particularly advantageous in practical settings, where data length can vary significantly due to different durations and measurement conditions.
Third, the method includes a learnable summary network that effectively reduces data dimensionality from potentially large time-series datasets to a fixed-size vector. Unlike traditional likelihood-based methods, where summary statistics are manually selected by practitioners, our method autonomously learns the most informative statistics from raw data.
Lastly, the proposed method offers significant computational advantages, particularly for problems that necessitate multiple running of the forward model from scratch given new datasets. This efficiency stems from the method’s implementation of amortized inference, which provides a clear benefit over ABC and traditional sampling-based methods. While the training phase has a relatively high computational cost, this phase can be conducted offline. Once trained, the networks estimate the posterior distributions using newly observed data in a matter of second.
The proposed method shows promise for real-time parameter estimation, yet future research should explore several promising directions. Currently, the method addresses only the inverse problem. However, developing a forward model or a likelihood emulator using normalizing flow is feasible, given its capacity for bijective mapping. Such an emulator could be integrated with various priors and MCMC methods to approximate posterior distributions.

### v2

@@ 3.5 Application
As described in Section 2.1, our case study focuses on estimating the hydraulic log-conductivity field of the 706 control volumes. The observed data consists of synthetic, noisy observations of the hydraulic head at 13 observation locations over 25 timesteps. The log-conductivity prior distribution is a Gaussian process prior with zero mean and the exponential covariance function. From this prior, 5,001 synthetic datasets are generated using MODFLOW. Fig. 7 illustrates a single realization of this synthetic data, showing the input conductivity field alongside the hydraulic head field across the simulation domain at the 25 timesteps. Cells colored white are inactive, meaning the governing equation is not solved in these cells, while the remaining cells are active. Additionally, to simulate measurement errors, random Gaussian noise with zero mean and standard deviation of 0.01 is added to the synthetic observations.
Figure 7: A realization synthetic data at 706 locations: (a) log-conductivity field; (b) hydraulic head field.
@@ 3.5.1 Model training
Fig. 8 shows the performance of the trained model for 100 test datasets. Due to space limitations in the paper, we show results for only a subset of the 706 parameters. The error bars depict the $95\%$ credibility intervals, calculated from 2,000 posterior samples. The results show that the estimated parameters generally align well with the ground truth, as evidenced by 87% (13 out of 15) of the selected parameters achieving a coefficient of determination $R^{2}$ above 0.8. Furthermore, the correlation coefficient $(r)$ confirms a strong linear relationship between the estimated posterior mean and actual values. However, it is noted that certain parameters, such as $\theta_{312}$ and $\theta_{315}$, exhibit lower $R^{2}$ values compared to others. This variation is attributed to their relative insensitivity to the observations collected from the 13 sensor locations. The 15 parameters shown were selected randomly for illustrative purposes.
Figure 8: Validation of trained model performance.
@@ 3.5.2 Results of parameter estimation
We compare our method against PEST-IES for two test problems corresponding to two reference conductivity fields. It is important to highlight the workflow differences between the methods: PEST-IES is likelihood-based and thus relies on a forward model to estimate parameters, whereas our proposed method bypasses the evaluation of the likelihood function. We employed three metrics to compare the accuracy of the parameter estimates derived from both methods:
@@
    where $\theta_{i,\text{est}}$ and $\theta_{i,\text{ref}}$ are the $i$th estimated and true parameters, respectively. A lower relative error signifies higher accuracy in the parameter estimation.
@@
    A higher LPP value points to a more accurate model estimation because it reflects a better fit of the estimated parameters to the true values.
Figs. 9 and 10 display the posterior mean log-conductivity fields estimated with PEST-IT and the likelihood-free method, true log-conductivity fields, and point relative errors for tests #1 and #2, respectively. It can be seen that both methods yield comparably accurate estimations, which suggests that the proposed method can reliably reconstruct the conductivity field without the need for likelihood function evaluation. Figs. 11 and 12 show the true parameter values, the posterior mean values, and the corresponding $95\%$ credibility intervals (shown in grey shaded areas) for both PEST-IES and the proposed method. A majority of the estimated parameters are contained within the credibility interval, which suggests that the proposed method is reliable and yields estimations that are comparable to those obtained via the PEST-IES benchmark method. Furthermore, these figures highlight that parameter estimation for test #2 poses a greater challenge, evidenced by the larger variability among the 706 parameters, particularly in the 200–300 and 500–600 location ranges, in comparison to test #1. Despite these challenges, the proposed method manages to produce accurate parameter estimations, demonstrating its efficacy even under challenging conditions.
Figure 9: Estimation of log-conductivity field using test #1: (a) PEST-IES, (b) proposed.
Figure 10: Estimation of log-conductivity field using test #2: (a) PEST-IES, (b) proposed.
Figure 11: Estimation of 706 parameters using test #1.
Figure 12: Estimation of 706 parameters using test #2.
Table 1 presents the accuracy metrics for the proposed PEST-IES methods for two test problems. For test #1, both methods exhibit comparable accuracy, as measured by the relative $\ell_{2}$ error and the LPP, with the proposed method showing a marginally lower average relative error compared to PEST-IES. When evaluating test #2, the proposed method demonstrates slightly better performance, with a lower average relative error, a lower relative $\ell_{2}$ error, and a marginally higher LPP score. While these differences are modest, they suggest that the proposed method maintains its precision in parameter estimation, even with the increased variability among the 706 parameters in test #2. Moreover, while the proposed and PEST-IES methods show similar performance levels in parameter estimation, the computational efficiency between the two is noteworthy (details are presented in Sec. LABEL:sec:cost). The PEST-IES method requires repeated evaluations of the forward model, which is computationally intensive. In contrast, the proposed method does not require the forward model evaluations outside the training data generation stage.
Table 1: Comparison between proposed and PEST-IES method.
@@
Fig. 13 shows the coverage map for each method and each of the two datasets under consideration. In this figure, yellow control volumes indicate that the true value is within the $95\%$ credibility interval, and a red control volume indicates that the true parameter is outside said interval. For test #1, PEST-IES and the proposed methods demonstrate similar coverages, indicating comparable reliability. However, for test #2, which presents a more complex challenge for parameter estimation, both methods yield lower coverage compared to test #1. Despite this, the proposed method achieves a higher coverage value closer to the expected value of $95\%$.
Figure 13: Parameter coverage: (a) test #1; (b) test #2.
Some likelihood-free approaches, such as ABC, and likelihood-based methods, including PEST-IES, require resampling or solving minimization problems for each new set of measurements. One of the attractive attributes of the proposed method is that it does not have to be retrained when measurements change. It can perform parameter estimation across different measurement set sizes utilizing a single trained model. To illustrate this adaptability, we considered five scenarios where the number of measurements at each location varies from 20 to 25 with a time increment of one year between measurements. In the training dataset, the number of observations at each location is 25. Fig. 14 shows the relative $\ell_{2}$ errors for various numbers of measurements. As anticipated, the relative $\ell_{2}$ errors for both test problems generally decrease as the number of observations increases. Importantly, the findings in Fig. 14 confirm that the proposed model, once trained, can estimate parameters from datasets of arbitrary sizes.
Figure 14: Parameter estimation by varied data length: (a) dataset #1; (b) dataset #2.
@@ 3.5.3 Observation prediction
In this section, we evaluate the posterior predictive distribution of hydraulic head measurements obtained using PEST-IES and our proposed method. The posterior predictive distribution is calculated by inputting the estimated log-conductivity field into the MODFLOW model. The proposed method is used to generate 2,000 posterior samples for each parameter, which are fed into the MODFLOW model, resulting in the same number of observation predictions. Figs. 15 and 16 display the predictive observation distribution obtained via PEST-IES and the proposed methods, with the grey shaded areas representing the corresponding $95\%$ credibility intervals. It can be seen that both methods are capable of accurately predicting observations across all locations and timesteps. For the PEST-IES method, the predictions closely coincide with the true values for both datasets with a notably small uncertainty, as indicated by the tight credibility intervals. In contrast, the proposed method exhibits larger uncertainties, and the credibility intervals span a wider range than those of PEST-IES, especially for observations at locations #1 through #4. This difference can be attributed to the fact that PEST-IES performs parameter estimation using a forward model that directly generates the synthetic observation data, thus bypassing the need for model training. Consequently, PEST-IES can yield more accurate predictions, albeit with the trade-off of requiring more computational resources due to the repeated evaluation of the forward model.
@@
In the proposed method, the errors due to the summary and inference network training contribute to the parameter estimation error. There are several error sources in the network models. Firstly, the optimization of the loss function inevitably introduces some error due to the finite number of training data, as described in Eq. 15. This error arises from the stochastic nature of the sampling process used during training. Secondly, the summary network choice might be suboptimal, potentially failing to capture all crucial information. Additionally, the designed invertible network might not sufficiently capture the complexities of the forward model’s behavior. For instance, the depth of the inference network in our study, which comprises 10 invertible blocks, may be inadequate for handling the 706-dimensional parameter space. Thirdly, fine-tuning hyperparameters such as the dropout rate, learning rate, and optimization method presents a complex challenge, with limited guidance available for the ideal design and training of neural networks. These errors can propagate into the predictive performance, affecting the uncertainty in observation predictions. Despite these larger uncertainties, the mean predictions from the proposed method consistently align well with the true observations and remain within the confidence intervals, underscoring the method’s reliability. Moreover, for test #2, both the uncertainty and deviation from the true observations are more pronounced than for test #1 because the proposed method finds it more challenging to estimate parameters for test #2—this is evidenced by the higher relative $\ell_{2}$ error reported in Table 1.
@@
Figure 15: Observation prediction at 13 locations for test #1: (a) proposed method; (b) PEST-IES.
@@
Figure 16: Observation prediction at 13 locations for test #2: (a) proposed method; (b) PEST-IES.
@@
The proposed method requires approximately 4.5 hours for training, but once this phase is complete, it can perform parameter estimation for a given measurement dataset in about 2 seconds. In contrast, PEST-IES requires around 2.5 hours to complete parameter estimation for each measurement dataset, as the IES algorithm must be restarted from scratch for each measurement dataset. Although the proposed method demands more time for training, this process can be conducted offline using simulated data. The amortized nature of the proposed approach allows for real-time inference and is valuable for parameter estimation problems with varying parameters and excitations that have to be estimated from new measurement data, such as varying boundary conditions to be estimated from time-series pressure observations.
@@ 4 Conclusion
In this work, we introduce a NF-based likelihood-free framework for tackling high-dimensional inverse problems. The proposed method includes three phases. The first phase involves generating synthetic training data from a forward model utilizing prior knowledge of the model parameters. In the second phase we train jointly the model’s inference and summary networks. The summary network autonomously learns the most informative representation of the observation data, ensuring that the features extracted are maximally informative for inference tasks. The inference network is designed using an architecture that combines cACL and SL layers. The summary features are passed to the inference network, allowing for the joint optimization of the two networks. The final phase consists of real-time inference. Once the model is trained, the posteriors of the model parameters can be computed in real time. We apply this framework to estimate the 706-dimensional hydraulic conductivity field using noisy and incomplete observations from the Freyberg groundwater model. Additionally, we conduct a comparative analysis between our method and the likelihood-based benchmark method, PEST-IES. The results highlight that our method can accurately and efficiently estimate the high-dimensional conductivity field. Below, we summarize the main advantages of our proposed method.
First, the proposed method represents a fully likelihood-free approach that directly estimates the posterior without the need to evaluate a likelihood function and does not rely on assumptions on the structure of the true posterior beyond the expressivity of the inference network. Second, the method exhibits robust performance across varied measurement sequences. In other words, it can handle variable dataset sizes with a single trained model, that is, without the need of re-training for different measurement sequences. This flexibility is particularly advantageous in practical settings when varying parameters must be estimated from new measurement data. Third, the method includes a learnable summary network that effectively reduces data dimensionality from potentially large time-series datasets to a fixed-size vector. Unlike traditional likelihood-based methods, where summary statistics are manually selected by practitioners, our method autonomously learns the most informative features from raw data. Lastly, the proposed method offers significant computational advantages, particularly for problems that necessitate multiple runs the forward model from scratch given new datasets. This efficiency stems from the method’s implementation of amortized inference, which provides a clear benefit over ABC and traditional sampling-based methods. While the training phase has a relatively high computational cost, this phase can be conducted offline. Once trained, the model can efficiently estimate the posterior distribution.